OpenStudy (anonymous):

1/sinx + 1/cosx=(cosx+sinx)(secx)(cscx) @zzr0ck3r

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

distribute the secx and cscx on the right and tell me what you have

OpenStudy (mathmale):

Cupcake: pretty please, include instructions when you post a problem. Have you tried adding the two fractions on the left AFTER finding and using the LCD?

OpenStudy (anonymous):

??@mathmale

OpenStudy (anonymous):

I need help with this...

OpenStudy (anonymous):

@zzr0ck3r Idk what I even did please help

OpenStudy (anonymous):

@zz0ck3r help

OpenStudy (zzr0ck3r):

you have (a+b)*c*d this is the same thing as acd+bcd I distributed the cd Do the same thing with sec(x)csc(x)

OpenStudy (zzr0ck3r):

what do you get?

OpenStudy (mathmale):

Starting with 1/sinx + 1/cosx=(cosx+sinx)(secx)(cscx) I'd write this a bit differently for simplicity: \[\frac{ 1 }{ \sin x }+\frac{ 1 }{ \cos x }\]

OpenStudy (mathmale):

Then I'd identify the LCD (lowest common denominator) of these two fractions. What is it? Combine these two fractions using the LCD. Please show your work so i can give you feedback on it.

OpenStudy (mathmale):

Come on!! To find the LCD, just multiply: (sin x)(cos x). Then re-write that sum of fractions as \[\frac{ \cos x }{ (\sin x)(\cos x) }+\frac{ \sin x }{ (\sin x)(\cos x) }.\] what next in the problem solution? (Hint: try looking up trig identities and applying the appropriate ones.)